The Moon has a mass of M = 7.3 · 1022 kg, a radius of R = 1.7 · 106 m and a rotation period of
T = 27.3 days. Scientists are planning to place a satellite around the Moon that always remains
above the same position (geostationary).
(a) Calculate the distance from the Moon’s surface to this satellite.
(b) Explain if such a Moon satellite is possible in reality.
1
Expert's answer
2019-05-13T10:49:19-0400

In the case of a stationary circular rotation of a satellite of mass m, the Moon’s gravitational force (gravity) Fg and the centrifugal force Fc act on it, they balance each other (Fg=Fc).


"Fc=mv^2\/(R+h)"

where v - speed of the spacecraft, m - mass of the spacecraft, R - radius of the moon, h - height of the spacecraft above the surface of the moon.


"Fg=GMm\/(R+h)^2"

G – gravitational coVnstant, M - the mass of the Moon

Equating forces and using the formula for speed


"v=2\pi(R+h)/T"

(T- rotation period):



"h=cubic root(T^2GM/(4\pi^2))-R"

"h=8.7*10^7 m"

"h>>R"

Too far from the moon, this is impossible